# Taxonomy-Guided Zero-Shot Recommendations with LLMs

### v2

@@ 4.1.4 Implementation Details
To ensure uniform sequence lengths, we use the user’s last interacted item to pad the historical interaction sequences with lengths < the threshold, 10. Because we are studying the ability to use LLM as a personal recommendation assistant, we focused on close-source and API-based LLMs in the evaluation to fit our scenario. Specifically, we evaluate the widely used GPT-4 by OpenAI’s API, and conduct each experiment three times and present the average results.
@@ 4.2 Overall Performance (RQ1)
In this section, we aim to investigate the recommendation performance of various methods under the zero-shot setting, which enables us to evaluate how LLMs can be used as recommenders without tuning parameters or any historical user-item interactions. The evaluation results are presented in Table 2. We compare our proposed TaxRec with two types of models: traditional pre-trained zero-shot recommendation models and LLM-based zero-shot models, as above the line and under the line respectively.
From the table, we can draw the following observations: (1) Our proposed TaxRec significantly outperforms both traditional and LLM-based methods, demonstrating the superiority of prompting LLM with our proposed taxonomy framework to make recommendations in the zero-shot scenario. TaxRec successfully retrieves LLM’s knowledge to facilitate the generation ability for recommendation task without any knowledge outside of LLMs. In this way, TaxRec successfully unifies the recommendation task to the NLP task. (2) LLM-based zero-shot method, i.e., DirectRec, has limited recommendation ability. For example, in the Movie dataset, DirectRec achieves comparable performance compared with traditional methods, however, in the Book dataset, it can hardly make correct recommendations. We can infer that in some domains that LLMs have seen before, such as Movie, their performance can be similar to traditional pre-trained models. While in some domains that LLMs have not seen before, such as Book, their recommendation capability is impeded. Nevertheless, after applying our proposed taxonomy method, LLMs can achieve significantly better performance, almost 10 times greater than DirectRec in Book dataset. These results show that there is still a gap between language task and recommendation task when using LLMs, which indicates the importance of our study. Additionally, it demonstrates that our taxonomy approach unlocks the potential of LLMs on recommendation tasks. (3) Some traditional pre-trained recommendation models can achieve a fair performance. For example, RecFormer and ZESRec perform well in Movie dataset, and Popularity performs the second best in Book dataset. This implies that each domain may have different characteristics and thus be suitable for different methods. However, our proposed TaxRec performs best for both datasets, showing that it has fully tapped into LLM’s profound knowledge.
@@ 4.3 Ablation and Effectiveness Analysis (RQ2)
This section conducts ablation experiments on TaxRec by removing taxonomy regularization and the feature-based matching separately. The results are shown in Table 3. "w/o Tax" uses LLMs to make recommendations given user’s original historical interactions without taxonomy. "w/o Match" is TaxRec without taxonomy-instructed matching mechanism, which directly maps LLM’s generated text to the original item pool.
Table 3: Ablation Study of TaxRec
@@
To investigate how taxonomy can help retrieve LLM’s knowledge, we conduct an experiment on TaxRec’s variant "w/o Tax", which uses LLMs to make recommendations without taxonomy (only with user’s original historical interactions). From this table, we can observe that the performance of "w/o Tax" is significantly lower than TaxRec. In the Movie dataset, "w/o Tax" can just achieve half of the ability of TaxRec, while in the Book dataset which LLMs may not have good knowledge of, we can see a 10-times drop in the performance of "w/o Tax" compared with TaxRec. These results show that our designed taxonomy method plays an essential role in LLM recommendation. With the help of taxonomy, LLM-embedded knowledge can be better retrieved, which makes great use of LLM’s capability for recommendation tasks.
Though taxonomy can help us retrieve LLM’s knowledge, the outputs of LLMs are still plain text. From Table 3, we can find that without our parsing and matching mechanism, i.e., "w/o Match", the recommendation performance will decrease for both datasets. Because the generated text is unstructured strings, directly calculating the similarity and mapping it to the candidate items will make LLMs confused and lose some information. However, after parsing the outputs to the taxonomy’s format and matching them with the categorized item pool in the same format, the recommendation performance is better. These results demonstrate that when paired with the taxonomy-instructed matching, TaxRec will have better performance, implying the effectiveness of the matching.
@@
4.4 Hyperarameter Analysis (RQ3)
In our proposed TaxRec, there are two hyperparameters that make important impacts on recommendation performance: 1) the number of features in the taxonomy, and 2) the methods for calculating the matching score.
Number of Features in Taxonomy. Since our proposed TaxRec improves LLM recommendation by using an intermediate taxonomy, where the taxonomy is a list of features, the number of features is key to performance. As the taxonomy is generated by LLMs, the total number of features may differ per domain (dataset). To investigate this hyperparameter’s impact, we vary the number of features for each dataset and present the results in Figure 3. We find that: (1) Generally, more features lead to better recommendation performance. For example, TaxRec with a 5-feature taxonomy yields the lowest results across both metrics, as more features enrich item representation and retrieve more domain knowledge from LLMs, enhancing recommendations. (2) However, more features do not always equate to better performance. In Figure 3(a), Recall@5 and Recall@10 slightly decrease when using 20 features compared to 10. Similarly, in Figure 3(b), NDCGs drop when using 15 features instead of 10. We infer that exceeding the LLMs’ domain knowledge leads to some features being assigned random values, introducing noise and reducing recommendation quality compared to using fewer, more relevant features.
@@ 4.4 Hyperarameter Analysis (RQ3)
Methods for Matching Score. The matching component is essential in our TaxRec, in which the method that calculates the matching score is the key. To investigate the effect of it, we examine two different types of methods, the learning-based method and the rule-based method. BERT embedding is a representative learning-based method that can be used to calculate the matching score. It leverages the contextual understanding of pre-trained models and can capture semantic similarities between texts. While BLUE score, ROUGE score, and our proposed taxonomy-instructed matching mechanism are rule-based methods. They rely on predefined rules and algorithms to evaluate the similarity, each with a different focus. Figure 4 shows the results of using these methods. From the figure, we can have the following findings: (1) The learning-based method, i.e., BERT embedding, does not perform well on both datasets. This is because the model is not pre-trained on our dataset, and the semantic information is not our focus. The learning-based method can perform well if they are well pre-trained or fine-tuned on a specific dataset, however, the training will cost some time and resources. (2) Instead, rule-based methods are more suitable for TaxRec. The reason is that TaxRec uses a taxonomy to structure both LLM’s outputs and the candidate item pool, thus the representations for them are mixed with fragmented information rather than semantic information. (3) Among the rule-based methods, our proposed matching mechanism performs best. This is because this mechanism is instructed by the taxonomy which specifically aligns with our task. While instructed by taxonomy, the outputs generated by LLMs can be easily parsed and the similarity can be easily calculated by word-wised text matching without other sophisticated rules.

### v3

@@ 4.1.4 Implementation Details
To ensure consistent sequence lengths, we pad historical interaction sequences shorter than the threshold (10) with the user’s most recent interaction. For the LLM evaluation, we use both closed-source (GPT-4 via OpenAI’s API) and open-source (LLaMA-2-7b with pre-trained parameters), which are widely adopted. Each experiment is repeated three times, and the average results are reported.
@@ 4.2 Overall Performance (RQ1)
In this section, we evaluate the recommendation performance of various methods in a zero-shot setting, which allows us to assess how LLMs can be utilized as recommenders without parameter tuning. The results are presented in Table 2. We compare our proposed TaxRec with two categories of models: traditional pre-trained zero-shot recommendation models (above the line) and LLM-based zero-shot models (below the line).
The following key observations can be drawn from the table: (1) Our proposed TaxRec significantly outperforms both traditional and LLM-based methods, particularly when applied to GPT-4, showcasing the effectiveness of prompting LLMs with our taxonomy framework in a zero-shot scenario. TaxRec leverages the LLM’s internal knowledge to facilitate recommendation generation without relying on external information, thus unifying the recommendation task with the NLP task. (2) The LLM-based zero-shot method DirectRec shows limited recommendation capability. For instance, while DirectRec performs comparably to traditional models on the Movie dataset, it struggles on the Book dataset, where it barely produces correct recommendations. This suggests that LLMs perform better in domains they have encountered before, like Movies, but face challenges in unfamiliar domains, such as Books. However, by applying our taxonomy framework, LLMs achieve substantially better performance—nearly ten times higher than DirectRec on the Book dataset. These findings highlight the gap between language tasks and recommendation tasks when using LLMs, reinforcing the importance of our study. Furthermore, it demonstrates how our taxonomy approach unlocks the potential of LLMs for recommendation tasks. (3) The performance improvements of TaxRec vary depending on the underlying LLM. For example, the improvement of TaxRec-GPT4 over DirectRec-GPT4 is more pronounced than the improvement of TaxRec-LLaMA2 over DirectRec-LLaMA2. This could be attributed to the inherent capabilities of different LLMs, such as comprehension and generation. Despite these differences, our proposed TaxRec consistently enhances the performance of direct recommendations by LLMs, underscoring its effectiveness regardless of the LLM used.
@@ 4.3 Ablation and Effectiveness Analysis (RQ2)
In this section, we conduct ablation studies on TaxRec to analyze the effectiveness of its component design and prompt design.
Component Design. TaxRec consists of two key components: taxonomy regularization and feature-based matching. We perform ablation experiments by separately removing each component, with results shown in Table 3. In the “w/o Tax” variant, LLMs generate recommendations based solely on the user’s original historical interactions, without the taxonomy. The “w/o Match” variant excludes the taxonomy-instructed matching mechanism and instead directly maps LLM-generated text to the original item pool.
@@
Table 3: Performance of different component design variants of TaxRec.
The results show that the “w/o Tax” variant performs significantly worse than TaxRec. In the Movie dataset, “w/o Tax” achieves only half of TaxRec’s performance, while in the Book dataset, where LLMs may have limited prior knowledge, performance drops by nearly tenfold compared to TaxRec. These results highlight the crucial role of taxonomy in LLM-based recommendation. The taxonomy helps retrieve the LLM’s internal knowledge more effectively, enhancing its ability to perform recommendation tasks.
Although taxonomy retrieval enhances LLM performance, the raw outputs from LLMs are still unstructured text. Table 3 shows that the absence of our parsing and matching mechanism (“w/o Match”) results in reduced performance in both datasets. Without structured parsing, direct similarity calculations and mappings to candidate items cause LLMs to lose important information. By parsing outputs into the taxonomy format and matching them with the categorized item pool, recommendation accuracy is significantly improved. This demonstrates that taxonomy-instructed matching further boosts TaxRec’s performance, underscoring its effectiveness.
Variant Movie Book h w/ t h w/o t h w/ t h w/o t rec w/ title 0.235 0.265 0.240 0.200 rec w/o title 0.300 0.180 0.225 0.025
Table 4: Prompt design variants of TaxRec. “h w/ t” refers to history with title, “h w/o t” to history without title; “rec w/ title” refers to recommendation with title, and “rec w/o title” to recommendation without title. Results are Recall@10.
Prompt Design. We also experimented with different prompt templates to optimize TaxRec. Our framework uses three key prompts: the Taxonomy Generation Prompt, the Categorization Prompt, and the Recommendation Prompt. Table 4 summarizes various prompt configurations we tested, such as including or excluding item titles in the historical sequences and recommendations.
The results show that the optimal prompt design can vary across datasets. For the Movie dataset, the best combination was representing the history with titles but generating recommendations without titles. These variations emphasize the importance of prompt design in achieving optimal performance.
@@
4.4 Hyperparameter Analysis (RQ3)
In TaxRec, two hyperparameters have a significant impact on recommendation performance: (1) the number of features in the taxonomy, and (2) the method for calculating the matching score.
Number of Features in Taxonomy. TaxRec improves LLM-based recommendations using an intermediate taxonomy of features, where the number of features is key to performance. Since the taxonomy is generated by LLMs, the feature count may vary across datasets. Varying the number of features, we observe that: (1) Generally, more features lead to better performance, as they provide richer item representation and leverage more domain knowledge from LLMs. For instance, using just 5 features results in the lowest performance across metrics. (2) However, too many features can reduce performance. In Figure 3(a), Recall@5 and Recall@10 drop slightly when moving from 10 to 20 features. Similarly, in Figure 3(b), NDCG declines with 15 features compared to 10. This suggests that excessive features may introduce noise by exceeding the LLM’s domain knowledge.
@@ 4.4 Hyperparameter Analysis (RQ3)
Methods for Matching Score. The matching component is essential in our TaxRec, in which the method that calculates the matching score is the key. We evaluate two types of methods: learning-based methods and rule-based methods. BERT embedding is a representative learning-based method, capturing semantic similarities using pre-trained models. In contrast, rule-based methods like BLUE, ROUGE, and our taxonomy-instructed mechanism use predefined rules to assess similarity.
Figure 4 presents the results: (1) The learning-based method, i.e., BERT embeddings, performs poorly on both datasets. This is likely because the model is not pre-trained on our specific dataset, and semantic similarity is not the focus of our task. While learning-based methods can excel when pre-trained or fine-tuned on a specific dataset, such training is resource-intensive and time-consuming. (2) In contrast, rule-based methods are better suited to TaxRec. Since TaxRec structures both the LLM outputs and the candidate item pool using a taxonomy, the representations are composed of fragmented rather than purely semantic information. (3) Among the rule-based methods, our proposed taxonomy-instructed matching mechanism performs best. This is because it aligns directly with the taxonomy, allowing LLM outputs to be easily parsed, and enabling similarity to be calculated through word-level matching without the need for complex rules.
@@ 6 Limitations
Acknowledgments
This material is based upon work supported by the U.S. Department of Homeland Security under Grant Award Number 17STQAC00001-07-04, NSF awards (SaTC-2241068, IIS-2339198, III-2106758, and POSE-2346158), a Cisco Research Award, and a Microsoft Accelerate Foundation Models Research Award. The views and conclusions contained in this document are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the U.S. Department of Homeland Security and the National Science Foundation.